OpenStudy (abdumu16):

Find the equation of the tangent line of the given function at the point with the indicated x-coordinate. f(x) = x + (1/x); x = 4 y= ? (find the equation)

OpenStudy (sshayer):

when x=4 y=?

OpenStudy (sshayer):

\[f \prime(x)=1-\frac{ 1 }{ x^2 }\] at x=4 \[f \prime \left( x \right)=1-\frac{ 1 }{ 16 }=\frac{ 15 }{ 16 }\] eq. of tangent is \[y-y1=f \prime \left( 4 \right)\left( x-4 \right)\] when x=4, \[y=4+\frac{ 1 }{ 4 }=\frac{ 17 }{ 4 }\] \[y1=\frac{ 17 }{ 4 }\]